User Tools

Site Tools


lbaops:lbamar2008:v252ecdlog

Setup:

RCP into Chan 1 on Frequency Translator, LCP into Chan 2

Sky Freq 8425, Agilent 13050, Fixed 4100 = IF 525, SMY LO 627

DAS profile = VSOP.PRO

used 10th harmonics for coherence at 841.4 (→8414) and 843.0 (→8430)

Recorded to disks ATNF V011B, labels ATNF V011 B

Tried to repair faulty RF cable in receiver room carrying LCP, but could not. Replaced with a shorter, but higher loss cable a short time after the experiment started. Since we were unable to do system temperature scans before this experiment with the new setup, we are using the calibration from the previous 8.4 GHz experiment this session. Due to this, the LCP was not connected until about 20min into the exp. We probably also missed the first minute or so of RCP.

Was suspicious about having a much higher noise level than Hobart, despite having a similar system temperature. It was thought that maybe the maser reference was faulty. We found the oscillator that was used to generate coherence tones was left powered on (tone off) without being locked on the external reference, which may have caused the site-wide reference to become unreliable. After switching the external reference back on on this oscillator, a fringe test revealed that the noise level returned to a comparable value to Hobart's, suggesting the problem was solved by this action.

lbaops/lbamar2008/v252ecdlog.txt · Last modified: 2015/12/18 16:38 (external edit)